OpenStudy (anonymous):

t^2-49/t^2+4t-21 * t^2+8t+15/t^2-2t-35 multiply or divide

OpenStudy (anonymous):

\[\frac{ t^2-49 }{ t^2+4t-21 } * \frac{ t^2+8t+15 }{ t^2-2t-35 }\]

OpenStudy (anonymous):

first you need to factor everything

OpenStudy (anonymous):

You know how to do that or do you need help factoring? ^^

OpenStudy (anonymous):

im ok at factoring

OpenStudy (anonymous):

see if you can factor: t^2-49 t^2+4t-21 t^2+8t+15 t^2-2t-35

OpenStudy (anonymous):

(t+7)(t-7)/(t+3)(t-7)

OpenStudy (anonymous):

Close. (t+3)(t-7) would be t^2-7t+3t-21 which would be t^2-4t-21 instead of +4t

OpenStudy (anonymous):

so the first fraction factors into\[\frac{ (t+7)(t-7) }{ (t+7)(t-3) }\]

OpenStudy (anonymous):

See if you can factor the other fraction

OpenStudy (anonymous):

(t+3)(t+5)/ i dont know im having trouble with the bottom

OpenStudy (anonymous):

Well, it has to be factors that multiply together to get -35, and add together to get -2 7 and 5 make 35, and because the 2 is negative the bigger factor also has to be negative. (t-7)(t+5)

OpenStudy (anonymous):

o okay i see .

OpenStudy (anonymous):

so you have: \[\frac{ (t+7)(t−7) }{ (t+7)(t−3) } * \frac{ (t+3)(t+5) }{ (t-7)(t+5) }\]

OpenStudy (anonymous):

Now you cross out any like terms

OpenStudy (anonymous):

would it be 0 or im doing t way wrong

OpenStudy (anonymous):

\[\frac{ (t-7)(t+3) }{ (t-3)(t-7) }\]

OpenStudy (anonymous):

So you're left with: \[\frac{ t+3 }{ t-3 }\]

OpenStudy (anonymous):

what happen to the (t-7) and (t+7)

OpenStudy (anonymous):

They were both negative so they cancled out

OpenStudy (anonymous):

canceled *

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

would t+3/t−3 be the answer or there another stoep

OpenStudy (anonymous):

step

OpenStudy (anonymous):

Thats it ^^

OpenStudy (anonymous):

okay thank you . i appericate the help

OpenStudy (anonymous):

Mhm^^